OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

#humans = a #horses = b a + b = 74 (all humans and horses have exactly one head) a * 2 + b * 4 = 196 (assuming all humans have 2 legs, and all horses have 4 legs) a = 74 - b \(\Rightarrow\) (74 - b)*2 + 4b = 196 \(\Rightarrow\) 148 -2b +4b = 196 \(\Rightarrow\) 2b = 48 \(\Rightarrow\) b = 24 \(\Rightarrow\) a = 74 - 24 = 50